sarah invested $1200 in the bank part at 9% per annum and the remainder at 4% per annum. after one year the 9% investment had $56 more interest. how much was invested at each rate?
Answer by stanbon(75887)

sarah invested $1200 in the bank part at 9% per annum and the remainder at 4% per annum. after one year the 9% investment had $56 more interest. how much was invested at each rate?
-------
0.09x - 0.04(1200-x) = 56
9x - 4*1200 +4x = 5600
13x = 5600+4800
13x = 10400
x = $800 (amt. invested at 9%)
---
1200 -x = 400 (amt. invested at 4%)
------------------
Cheers,
Stan H.
================